OpenStudy (anonymous):

If x and y are integers, and 3x + 2y = 13, which of the following could be the value of y ? A. 0 B. 1 C. 2 D. 3 E. 4

OpenStudy (shamil98):

3x + 2y = 13 3x + 2(2) = 13 3x = 9 x= 3 (3,2) so C. it's asking for integers which is why this is true.

OpenStudy (squirrels):

i used to have a six pack like that